# Joint Pedestrian and Vehicle Traffic Optimization in Urban Environments using Reinforcement Learning

### v1

@@ Abstract
Reinforcement learning (RL) holds significant promise for adaptive traffic signal control. While existing RL-based methods demonstrate effectiveness in reducing vehicular congestion, their predominant focus on vehicle-centric optimization leaves pedestrian mobility needs and safety challenges unaddressed. In this paper, we present a deep RL framework for adaptive control of eight traffic signals along a real-world urban corridor, jointly optimizing both pedestrian and vehicular efficiency. Our single-agent policy is trained using real-world pedestrian and vehicle demand data derived from Wi-Fi logs and video analysis. The results demonstrate significant performance improvements over traditional fixed-time signals, reducing average wait times per pedestrian and per vehicle by up to $\mathbf{67\%}$ and $\mathbf{52\%}$ respectively, while simultaneously decreasing total accumulated wait times for both groups by up to $\mathbf{67\%}$ and $\mathbf{53\%}$. Additionally, our results demonstrate generalization capabilities across varying traffic demands, including conditions entirely unseen during training, validating RL’s potential for developing transportation systems that serve all road users.
@@ I Introduction
Figure 1: The Craver Road corridor with an intersection (INT) that contains one primary traffic signal and four signalized crosswalks, along with seven midblock signalized crosswalks (MB1111–MB7777) that control pedestrian–vehicle interactions.
@@
Traffic congestion has become a silent tax on modern civilization. Each year, drivers in U.S. cities waste an average of $54$ hours stuck in traffic [1], costing over $160$ billion in lost productivity [2]. With urban cores in metropolitan areas experiencing an increase of traffic inflow—up to $+25\%$ year-over-year [3]—congestion is set to worsen as urbanization continues. To address this challenge, traffic control systems have evolved from traditional handcrafted rules and actuated systems to Adaptive Traffic Signal Control (ATSC). Driven by cost-effectiveness [4, 5], increasing availability of traffic data [6], and advances in optimization techniques [7], ATSC has become a central focus of intelligent transportation systems research [8]. While ATSC has reduced congestion and improved vehicular flow, its evolution has largely overlooked a critical stakeholder: pedestrians.
@@
At the same time, pedestrian fatalities in the U.S. have reached their highest level in $41$ years, averaging about $21$ deaths per day [9, 10]. Urban areas are particularly affected, with $84$% of these fatalities occurring in cities and $76$% taking place at non-intersection locations such as mid-block crossings and commercial strips [11]. A key factor is the prevalence of unsignalized mid-block crosswalks, where pedestrians lack clear guidance and drivers may not yield, increasing accident risks [12]. Implementing signalized crosswalks at these locations—with dedicated pedestrian phases and clear right-of-way signals—has proven effective in reducing vehicle-pedestrian conflicts and enhancing rule compliance [13]. The inclusion of pedestrian-friendly features in ATSC frameworks ensures that traffic control systems not only alleviate congestion but also proactively protect vulnerable road users.
@@
As urban traffic control grows increasingly complex, ATSC strategies are leveraging real-time sensor data and computational methods. RL, in particular, has emerged as a leading approach due to its ability to learn adaptive policies directly from data without relying on traffic models or handcrafted rules. Recent methods have primarily focused on improving vehicular throughput, often neglecting pedestrian efficiency or modeling pedestrian behavior in overly simplified scenarios [14, 15, 16]. This vehicle-centric approach fails to capture the complex interactions between vehicles and pedestrians, especially in urban environments characterized by high pedestrian volumes, frequent mid-block crossings, and diverse mobility demands. Developing control policies that simultaneously optimize pedestrian and vehicular efficiency in realistic urban settings remains an open challenge.
@@
This paper introduces a deep RL framework for corridor-level control of eight traffic signals, jointly optimizing for vehicular and pedestrian efficiency. Our contributions are:
@@
  • We develop a single policy that effectively manages high traffic volumes (up to $6,000$ pedestrians/hr and $558$ vehicles/hr) based on real-world demand data derived from Wi-Fi logs and video analysis.
@@
  • Our results demonstrate substantial performance improvements over traditional fixed-time signals, reducing average wait time per vehicle by up to $52\%$ and average wait time per pedestrian by up to $67\%$.
@@
  • We provide insights into the behaviors learned by our policy, including coordination of multiple signals to create a “green wave” effect, and responsiveness to real-time traffic by switching phases more frequently.
@@
The code, data, and video demonstrations are available in GitHub repository: github.com/poudel-bibek/Urban-Control.
@@ II Related Work
Conventional adaptive traffic control systems have evolved through two primary approaches: model-based methods like SCOOT [17], which leverage macroscopic traffic flow equations to optimize signal timings, and rule-based systems such as SCATS [18], which select from a library of timing plans in response to detected traffic. Despite their widespread deployment and clear advantages over fixed-time control, these systems struggle to capture the inherent complexity and stochasticity of urban traffic. Common challenges include unpredictable variations in traffic volume and queue propagation through multiple intersections in the network [19]. To address these challenges, modern traffic control strategies harness real-time sensor data alongside advanced computational techniques—including machine learning [20, 21] and deep reinforcement learning (DRL) [22, 23]—effectively eliminating the need for explicit rules or pre-defined models.
@@
In particular, DRL has emerged at the forefront of modern traffic control research, demonstrating its effectiveness in both single-agent and multi-agent settings [24, 25]. In single-agent settings, DRL has been applied to optimize isolated intersections by selecting appropriate signal phases [26] or duration [27] based on observed traffic conditions, while multi-agent scenarios explore approaches where agents either collaborate to coordinate signals across networks [28] or compete to prioritize movement in traffic signal with higher demand [15]. These methods generally show better performance compared to fixed-time and conventional control approaches [29], demonstrating, for instance, reductions in average travel time and queue lengths [30]. Yet, it is worth noting that majority of existing studies remain primarily focused on optimizing vehicular throughput [16].
@@
However, real-world urban environments include both vehicles and pedestrians. Integrating pedestrian dynamics into ATSC systems introduces several complex challenges—such as ensuring pedestrian safety at crosswalks [14, 31] and balancing the needs of both vehicles and pedestrians to prevent excessive delays for one group while prioritizing the other [32]. Recognizing these complexities, recent studies have extended RL-based traffic signal control to incorporate pedestrians. Several approaches have emerged in this direction: some introduce pedestrian-specific phases i.e., eliminate vehicle-pedestrian interactions for improved safety [33], while others incorporate pedestrian-centric performance metrics directly into the reward function [34]. However, these studies typically rely on synthetic demand data or remain constrained to ideal road-networks [8]. This leaves pedestrian-inclusive ATSC strategies relatively underexplored, particularly for complex urban corridors with multiple signalized crosswalks.
@@
Our work addresses these limitations by proposing a DRL framework that jointly optimizes vehicular and pedestrian waiting times in a real-world corridor-level setting. Unlike earlier methods that either focused solely on vehicles or used synthetic demand on ideal networks, our approach employs a real-world urban network of eight traffic signals with real-world demand data for both vehicles and pedestrians. The work most similar to ours is Kumarasamy et al. [35], which also uses a real-world corridor and vehicle demand data but differs by adopting a multi-agent framework and relying on synthetic pedestrian demand. Our approach demonstrates that a single-agent policy can effectively control an entire urban corridor while balancing the needs of both vehicles and pedestrians.
@@ III Methodology
Figure 2: (a) Deep reinforcement learning framework for corridor-level traffic control. (b) Intersection signal configurations controlling vehicle and pedestrian movements, including dedicated left turns (choice $3$) and all-pedestrian phases (choice $4$). (c) Two midblock signal configurations allowing either vehicle movement (choice $1$) or pedestrian crossing (choice $2$).
III-A Real-World Network and Demand Data
The Craver Road, as shown in Figure 1, is a $750$m corridor that serves as the primary arterial through the University of North Carolina, Charlotte’s main campus. The campus covers $1.56$ square miles and includes $85$ buildings and approximately $34,000$ students, faculty, and staff [36]. To capture pedestrian demand data, we utilize Wi-Fi logs collected in September $2021$ by the university’s IT department. The Wi-Fi network comprises of $2,492$ access points (APs) distributed across $82$ buildings, with $88,409$ unique clients. The Wi-Fi log data captures communication events between Wi-Fi clients (e.g., smartphones and laptops) and APs. Each log entry indicates “when” (timestamp) and “where” (AP location) each client connected to the network. We processed this raw data using a generalized Wi-Fi processing framework based on a “Point-Line-Plane” hierarchical concept [37] with the following filtering assumptions:
  • Client activities were aggregated at the building level, each client’s AP sessions within the same building were merged to represent presence in that location.
@@ III-A Real-World Network and Demand Data
For vehicle demand data, we analyzed four video recordings (total $21$ minutes) captured at different times at the Craver Road intersection (INT). The observed flow was converted to hourly rates, resulting in an average headway of $18$ seconds between vehicles. We mapped both pedestrian and vehicle demand to SUMO [39] trip definitions, creating a network-wide real-world demand of $2223$ pedestrians/hr and $202$ vehicles/hr. For trip generation, we defined origin-destination pairs using Traffic Analysis Zones (TAZs). Pedestrian origin-destination pairs were derived from Wi-Fi building visit data, and vehicle pairs from movement records. In each trip, start and end points were assigned to specific edges within TAZs, with departure times based on observed timestamps.
@@ III-B Markov Decision Process
State: Traffic representation significantly impacts ATSC performance [40]. Our observation fuses vehicle and pedestrian occupancy data in areas neighboring each traffic signal, vehicles are detected within a 15151515100100100100m vicinity while pedestrians are detected within 555510101010m. To capture both spatial and temporal dynamics, we stack occupancy information over the action duration (each action step encompasses $N$ simulation timesteps). At action step $t$, the observation $o_{t}\in\Omega$ is formed by stacking $N$ vectors from the previous action interval:
@@
    where $\mu_{i}^{mb}(o_{t},\theta)$ is the probability for signal $i$ and $b=1$ indicates permission for vehicle flow i.e., Midblock choice $1$ shown in Figure 2 (c).
The overall action $a_{t}=\text{concat}(a_{t}^{int},a_{t,1}^{mb},\ldots,a_{t,7}^{mb})$ is an $8$-digit string. Restricting the intersection action to pre-defined safe configurations inherently enforces safety (preventing conflicting pedestrian–vehicle greens) and reduces the exploration space for more sample-efficient learning. For additional safety, a $4$-timestep mandatory yellow phase is automatically introduced in all the signals via an internal mechanism before switching from a green to a red signal. Note that while direct control over the yellow phase duration is not permitted, the duration of any other phase can be controlled by repeatedly selecting the same action. Each action lasts for $10$ simulation steps ($N=10$).
Reward: To fulfill the multi-objective goal of minimizing both vehicle and pedestrian waiting times, we build on the success of the Maximum Wait Aggregated Queue (MWAQ) [41], and propose the Exponentially Increasing Maximum Wait Aggregated Queue (EI-MWAQ). While MWAQ uses the product of queue length and maximum waiting time to approximate the worst-case aggregate delay, EI-MWAQ normalizes those values and applies an exponential penalty. This strongly disincentivizes large queues and long waits. For the intersection, we compute:
@@
where $N^{\mathrm{int}}_{\mathrm{veh}}$ and $N^{\mathrm{int}}_{\mathrm{ped}}$ denote the counts of waiting vehicles and pedestrians at the intersection, $W^{\mathrm{int}}_{\mathrm{veh}}$ and $W^{\mathrm{int}}_{\mathrm{ped}}$ their respective maximum waiting times, and $|D|$ the number of incoming directions. For each midblock signal $i$, we compute:
@@
where $|D_{\mathrm{mb}}|$ is the number of incoming directions. These per-signal values are aggregated across all $7$ midblock signals using the $L_{2}$ norm:
@@
Figure 3: Performance comparison between three traffic control approaches across varying demand scales. The figure shows wait times for both pedestrians (left) and vehicles (right), measured as average wait time per pedestrian/vehicle (top) and total wait time for pedestrians/vehicles (bottom). Our RL agent consistently outperforms traffic signal control (Signalized) across all demand levels, reducing average vehicle wait times by up to $52\%$ (from $42.6$ to $20.5$ seconds per vehicle at $2$× demand) while decreasing pedestrian wait times by up to $67\%$ (from $6.0$ to $2.0$ seconds per pedestrian at $2$× demand). Despite providing protected crossing phases through signalization, our RL approach achieves pedestrian wait times comparable to or slightly better than unsignalizing mid-block crosswalks (Unsignalized), while delivering approximately $20\%$ lower average vehicle wait times at higher demands. The gray-shaded areas to the left and right indicate demand levels unseen during training ($<1$× and $>2.25$×), where our approach generalizes effectively with consistent performance improvements across both low and high demands. All values are averaged across $10$ independent simulation runs with a total of $600$ runs.
@@ IV Experiments
Simulation Time Step ($\Delta t$) $1$ second

### v2

@@ Abstract
Reinforcement learning (RL) holds significant promise for adaptive traffic signal control. While existing RL-based methods demonstrate effectiveness in reducing vehicular congestion, their predominant focus on vehicle-centric optimization leaves pedestrian mobility needs and safety challenges unaddressed. In this paper, we present a deep RL framework for adaptive control of eight traffic signals along a real-world urban corridor, jointly optimizing both pedestrian and vehicular efficiency. Our single-agent policy is trained using real-world pedestrian and vehicle demand data derived from Wi-Fi logs and video analysis. The results demonstrate significant performance improvements over traditional fixed-time signals, reducing average wait times per pedestrian and per vehicle by up to $\mathbf{67\%}$ and $\mathbf{52\%}$ respectively, while simultaneously decreasing total wait times for both groups by up to $\mathbf{67\%}$ and $\mathbf{53\%}$. Additionally, our results demonstrate generalization capabilities across varying traffic demands, including conditions entirely unseen during training, validating RL’s potential for developing transportation systems that serve all road users.
@@ I Introduction
Figure 1: The Craver Road corridor with an intersection (INT) that contains one primary traffic signal and four signalized crosswalks, along with seven mid-block signalized crosswalks (MB1111–MB7777) that control pedestrian–vehicle interactions.
@@
Traffic congestion has become a silent tax on modern civilization. Each year, drivers in U.S. cities waste an average of $54$ hours stuck in traffic [1], costing over $160$ billion in lost productivity [2]. With urban cores in metropolitan areas experiencing an increase of traffic inflow, up to $25\%$ year-over-year [3], congestion is set to worsen as urbanization continues. To address this challenge, traffic control systems have evolved from traditional handcrafted rules and actuated systems to Adaptive Traffic Signal Control (ATSC). Driven by cost-effectiveness [4, 5], increasing availability of traffic data [6], and advances in optimization techniques [7], ATSC has become a central focus of intelligent transportation systems research [8]. While ATSC has reduced congestion and improved vehicular flow, its evolution has largely overlooked a critical stakeholder: pedestrians.
@@
Currently, pedestrian fatalities in the U.S. have reached their highest level in $41$ years, averaging about $21$ deaths per day [9, 10]. Urban areas are particularly affected, with $84$% of these fatalities occurring in cities and $76$% taking place at non-intersection locations such as mid-block crossings and commercial strips [11]. A key factor is the prevalence of unsignalized mid-block crosswalks, where pedestrians lack clear guidance and drivers may not yield, increasing accident risks [12]. Implementing signalized crosswalks at these locations, with dedicated pedestrian phases and clear right-of-way signals, has proven effective in reducing vehicle-pedestrian conflicts and enhancing rule compliance [13]. The inclusion of pedestrian-friendly features in ATSC frameworks ensures that traffic control systems not only alleviate congestion but also proactively protect vulnerable road users.
@@
As urban traffic control grows increasingly complex, ATSC strategies are leveraging real-time sensor data and computational methods. Reinforcement learning (RL) has emerged as a leading approach due to its ability to learn adaptive policies directly from data without relying on traffic models or handcrafted rules. Recent RL-based methods have primarily focused on improving vehicular throughput, neglecting pedestrian efficiency or modeling pedestrian behavior in overly simplified scenarios [14, 15, 16]. This vehicle-centric approach fails to capture the complex interactions between vehicles and pedestrians, especially in urban environments characterized by high pedestrian volumes, frequent mid-block crossings, and diverse mobility demands. Developing control policies that simultaneously optimize pedestrian and vehicular efficiency in realistic urban settings remains an open challenge.
@@
We introduce an RL framework for corridor-level control of eight traffic signals, jointly optimizing for vehicular and pedestrian efficiency. Our contributions are
@@
  • we develop a single policy that effectively manages high traffic volumes (up to $6,000$ pedestrians/hr and $558$ vehicles/hr) based on real-world demand data derived from Wi-Fi logs and video analysis;
@@
  • our results demonstrate substantial performance improvements over traditional fixed-time signals, reducing average wait time per vehicle by up to $52\%$ and average wait time per pedestrian by up to $67\%$;
@@
  • we provide insights into the behaviors learned by our policy, including its ability to coordinate multiple independent signals to create a “green wave” effect and its responsiveness to real-time traffic, as demonstrated by its adaptive phase switching.
@@
The code, data, and videos are available in our GitHub: github.com/poudel-bibek/Urban-Control.
@@ II Related Work
Conventional adaptive traffic control systems have relied on model-based [17] or rule-based [18] approaches. Despite their widespread deployment and clear advantages over fixed-time control, these systems struggle to capture the inherent complexity and stochasticity of urban traffic. Common challenges include unpredictable variations in traffic volume and queue propagation through multiple intersections [19]. In response to these limitations, computational techniques that learn adaptive policies without relying on explicit rules or models, particularly machine learning and deep reinforcement learning (DRL), offer promising alternatives for traffic signal control [20, 21, 22, 23, 24, 25]. In single-agent settings, DRL has been applied to optimize intersection signal timing by selecting appropriate phase [26] or its duration [27] based on observed traffic conditions. While multi-agent scenarios explore approaches where agents either collaborate to coordinate signals across a network [28] or compete to prioritize movement at a traffic signal with higher demand [15]. These methods demonstrate improved performance over fixed-time and conventional approaches [29], achieving, for instance, reductions in average travel time and queue lengths [30]. Yet, it is worth noting that the majority of existing studies remain primarily focused on optimizing vehicle-centric outcomes such as reducing queue, decreasing stop frequency, and increasing throughput [16].
@@
However, real-world urban environments include both vehicles and pedestrians. Integrating pedestrian dynamics into ATSC systems introduces several complex challenges such as ensuring pedestrian safety at crosswalks [14, 31] and balancing the needs of both vehicles and pedestrians to prevent excessive delays for one group while prioritizing the other [32]. Recognizing these complexities, recent studies have extended RL-based traffic signal control to incorporate pedestrians. Several approaches have emerged in this direction: some introduce pedestrian-specific phases, i.e., eliminate vehicle-pedestrian interactions for improved safety [33], while others incorporate pedestrian-centric performance metrics directly into the reward function [34]. However, these studies typically rely on synthetic demand data or remain constrained to ideal road networks [8]. This leaves pedestrian-inclusive ATSC strategies relatively underexplored, particularly for complex urban corridors with multiple signalized crosswalks.
@@
Our work addresses these limitations by proposing a DRL framework that jointly optimizes vehicular and pedestrian waiting times in a real-world corridor-level setting. Unlike earlier methods that either focused solely on vehicles or used synthetic demand on ideal networks, our approach employs a real-world urban network of eight traffic signals with real-world demand data for both vehicles and pedestrians. The work most similar to ours uses a real-world corridor and vehicle demand data but differs by adopting a multi-agent framework and relying on synthetic pedestrian demand [35]. Our approach demonstrates that a single-agent policy can effectively control an entire urban corridor while balancing the needs of both vehicles and pedestrians.
@@ III Methodology
Figure 2: (a) Deep reinforcement learning framework for corridor-level traffic control. (b) Intersection signal configurations controlling vehicle and pedestrian movements, including dedicated left turns (choice $3$) and all-pedestrian phases (choice $4$). (c) Two mid-block signal configurations allowing either vehicle movement (choice $1$) or pedestrian crossing (choice $2$).
III-A Real-World Network and Traffic Data
Craver Road, shown in Figure 1, is a $750$m corridor that serves as the primary arterial through the University of North Carolina, Charlotte’s main campus. The campus covers $1.56$ square miles and includes $85$ buildings and approximately $34,000$ students, faculty, and staff [36]. To capture pedestrian demand data, we utilize Wi-Fi logs collected in September $2021$ by the university’s IT department. The Wi-Fi network comprises $2,492$ access points (APs) distributed across $82$ buildings, with $88,409$ unique clients. The Wi-Fi log data captures communication events between Wi-Fi clients (e.g., smartphones and laptops) and APs. Each log entry indicates “when” (timestamp) and “where” (AP location) each client connected to the network. We processed this raw data using a generalized Wi-Fi processing framework based on a “Point-Line-Plane” hierarchical concept [37] with the following filtering assumptions:
  • Client activities were aggregated at the building level; each client’s AP sessions within the same building were merged to represent presence in that location.
@@ III-A Real-World Network and Traffic Data
For vehicle demand data, we analyzed four video recordings (total $21$ minutes) captured at different times at the Craver Road intersection (INT). The observed flow was converted to hourly rates, resulting in an average headway of $18$ seconds between vehicles. We mapped both pedestrian and vehicle demand to SUMO [39] trip definitions, creating a network-wide real-world demand of $2223$ pedestrians/hr and $202$ vehicles/hr. Of these pedestrians, $990$ ($44$%) have trips that cross the corridor. For trip generation, we defined origin-destination pairs using Traffic Analysis Zones (TAZs). Pedestrian origin-destination pairs were derived from Wi-Fi building visit data, and vehicle pairs from movement records. In each trip, start and end points were assigned to specific edges within TAZs, with departure times based on observed timestamps.
@@ III-B Markov Decision Process
State: Traffic representation significantly impacts ATSC performance [40]. Our observation fuses vehicle and pedestrian occupancy data in areas neighboring each traffic signal. Vehicles are detected within a 15151515100100100100m vicinity while pedestrians are detected within 555510101010m. To capture both spatial and temporal dynamics, we stack occupancy information over the action duration (each action step encompasses $N$ simulation timesteps). At action step $t$, the observation $o_{t}\in\Omega$ is formed by stacking $N$ vectors from the previous action interval:
@@
    where $\mu_{i}^{mb}(o_{t},\theta)$ is the probability for signal $i$ and $b=1$ indicates permission for vehicle flow, i.e., Mid-block choice $1$ shown in Figure 2 (c).
The overall action $a_{t}=\text{concat}(a_{t}^{int},a_{t,1}^{mb},\ldots,a_{t,7}^{mb})$ is an $8$-component vector. Restricting the intersection action to pre-defined safe configurations inherently enforces safety (preventing conflicting pedestrian–vehicle greens) and reduces the exploration space for more sample-efficient learning. For additional safety, a $4$-timestep mandatory yellow phase is automatically introduced for all signals via an internal mechanism before switching from a green to a red signal. Note that while direct control over the yellow phase duration is not permitted, the duration of any other phase can be controlled by repeatedly selecting the same action. Each action lasts for $10$ simulation steps ($N=10$).
Reward: To simultaneously minimize wait times for vehicles and pedestrians, we propose the Exponentially Increasing Maximum Wait Aggregated Queue (EI-MWAQ) reward function. This reward is designed to reflect real-world behavior, where very long wait times lead to disproportionately high driver and pedestrian frustration. It builds on the Maximum Wait Aggregated Queue (MWAQ) [41], which uses the product of queue length and maximum waiting time to approximate worst-case delay, but introduces two modifications. First, penalties from all mid-block crossings are aggregated using an $L_{2}$-norm. Second, the final penalty values are obtained by normalizing these aggregate delays and applying an exponential function. This results in a penalty that grows rapidly as queue lengths and wait times increase. For the intersection, we compute:
@@
where $N^{\mathrm{int}}_{\mathrm{veh}}$ and $N^{\mathrm{int}}_{\mathrm{ped}}$ denote the counts of waiting vehicles and pedestrians at the intersection, $W^{\mathrm{int}}_{\mathrm{veh}}$ and $W^{\mathrm{int}}_{\mathrm{ped}}$ their respective maximum waiting times, and $|D|$ the number of incoming directions. For each mid-block signal $i$, we compute:
@@
where $|D_{\mathrm{mb}}|$ is the number of incoming directions. These per-signal values are aggregated across all $7$ mid-block signals using the $L_{2}$ norm:
@@ IV Experiments
Simulation Crossing Width $4$ m
Sidewalk Width $4$ m
Time Step ($\Delta t$) $1$ second
@@
Vehicle Control Model IDM [44]
Pedestrian Control Model Stripping [45]
Vehicle Speed Limit $50$ km/hr
Pedestrian Walking Speed $2.78$ m/s
@@ IV-A Setup
Figure 3: Performance comparison between three traffic control approaches at various demands. The figure shows wait times for both pedestrians (left) and vehicles (right), measured as average wait time per pedestrian/vehicle (top) and total wait time for pedestrians/vehicles (bottom). Our RL agent consistently outperforms the fixed-time signal control (Signalized) across all demand levels, reducing average vehicle wait times by up to $52\%$ (from $42.6$ to $20.5$ seconds per vehicle at $2$× demand) while decreasing pedestrian wait times by up to $67\%$ (from $6.0$ to $2.0$ seconds per pedestrian at $2$× demand). Despite providing protected crossing phases through signalization, our RL approach achieves pedestrian wait times comparable to or slightly better than unsignalizing mid-block crosswalks (Unsignalized), while delivering approximately $20\%$ lower average vehicle wait times at higher demands. The gray-shaded areas to the left and right indicate demand levels unseen during training ($<1$× and $>2.25$×), where our approach generalizes effectively with consistent performance improvements across both low and high demands. All values are averaged across $10$ independent simulation runs with a total of $600$ runs.